# SubZero: Random Subspace Zeroth-Order Optimization for Memory-Efficient LLM Fine-Tuning

### v1

@@ Abstract
Fine-tuning Large Language Models (LLMs) has proven effective for a variety of downstream tasks. However, as LLMs grow in size, the memory demands for backpropagation become increasingly prohibitive. Zeroth-order (ZO) optimization methods offer a memory-efficient alternative by using forward passes to estimate gradients, but the variance of gradient estimates typically scales linearly with the model’s parameter dimension—a significant issue for LLMs. In this paper, we propose the random Subspace Zeroth-order (SubZero) optimization to address the challenges posed by LLMs’ high dimensionality. We introduce a low-rank perturbation tailored for LLMs that significantly reduces memory consumption while improving training performance. Additionally, we prove that our gradient estimation closely approximates the backpropagation gradient, exhibits lower variance than traditional ZO methods, and ensures convergence when combined with SGD. Experimental results show that SubZero enhances fine-tuning performance and achieves faster convergence compared to standard ZO approaches like MeZO across various language modeling tasks.
22footnotetext: Corresponding author (email: jiali@bnu.edu.cn).33footnotetext: Our code is available at: https://github.com/zimingyy/SubZero.
@@ 1 Introduction
Large Language Models (LLMs), such as the GPT and LLaMA series (Zhang et al., 2022; Touvron et al., 2023), have recently demonstrated impressive capabilities in natural language processing tasks and beyond (Solaiman et al., 2019; Achiam et al., 2023). These models utilize deep learning, particularly the transformer architecture (Vaswani et al., 2017), to learn complex patterns in language data. However, LLMs can struggle with specialized tasks that require domain-specific knowledge (Shen et al., 2024). Fine-tuning presents an effective solution by slightly adjusting pre-trained LLMs with domain data, enabling them to adapt to specific tasks more effectively.
For fine-tuning, first-order (FO) optimizers, such as SGD (Amari, 1993) or Adam (Kingma & Ba, 2015), are commonly used to achieve promising performance on domain datasets. However, as LLMs grow in size, FO optimizers demand increasingly memory consumption due to the gradient computations required by backpropagation (BP) (Zhao et al., 2024a). To enhance memory efficiency, MeZO (Malladi et al., 2023) first introduces the zeroth-order (ZO) optimizer to LLM fine-tuning without BP. It just needs forward passes and calculates gradient estimates using finite differences of training loss values. Nevertheless, the variance of ZO gradient estimates linearly depends on the perturbation dimension, which corresponds to the number of model parameters. This can become extremely large in LLMs, resulting in significant performance degradation compared to FO optimizers (Gautam et al., 2024; Jiang et al., 2024; Liu et al., 2024).
There are two main attempts to addressing the high variance of ZO gradient estimates. The first approach involves increasing batch size alongside training iterations, which reduces gradient noise and variance in ZO gradient estimates (Gautam et al., 2024; Jiang et al., 2024). However, this leads to significant runtime and memory costs due to the large batch size in the later training stages. The second approach focuses on perturbing fewer parameters by employing sparse parameter perturbations, such as random and sparse pruning masks (Liu et al., 2024) and block-coordinate perturbations (Zhang et al., 2024), or by reducing the number of trainable parameters through techniques like parameter-efficient fine-tuning (PEFT) (Malladi et al., 2023; Zhang et al., 2024) and tensorized adapters (Yang et al., 2024). Recent theoretical advancements have proposed using random projections to lessen the dimensionality dependence in ZO optimizers (Nozawa et al., 2024; Roberts & Royer, 2023; Kozak et al., 2021) by applying low-dimensional perturbations in random subspaces. Nonetheless, a major drawback of this approach is the need to store a huge projection matrix that scales with model parameter dimensionality, making it impractical for fine-tuning large LLMs.
@@
Figure 1: Visualization of cosine similarity $\mathbb{E}\left[{\texttt{cosine}(\bm{g},\hat{\bm{g}})}\right]$, relative variance $\mathrm{Var}\bigl{[}\left\|\hat{\bm{g}}\right\|\bigr{]}/\left\|\bm{g}\right\|^{2}$, training loss, and GPU memory cost on OPT-1.3B under the prompt tuning scheme. Here, $\hat{\bm{g}}$ represents the gradient estimated by MeZO or our SubZero, and $\bm{g}$ denotes the expected gradient $\mathbb{E}[\hat{\bm{g}}]$. SubZero demonstrates reduced angle error and variance in gradient estimation, while also accelerating convergence with minimal additional memory overhead.
@@
Firstly, we propose a layer-wise low-rank perturbation approach for gradient estimation, specifically designed for fine-tuning LLMs. In each layer, we generate a low-rank perturbation matrix by combining two column-orthogonal matrices with a Gaussian random matrix, which is then used for gradient estimation. Unlike traditional ZO methods like MeZO (Malladi et al., 2023) which apply non-low-rank perturbations to the entire model, our approach significantly reduces the variance of gradient estimates and the angle error between the estimated gradient and its expectation, as respectively shown in Fig. 1 (a) and (b). SubZero also improves upon random subspace ZO methods like S-RGF (Nozawa et al., 2024) by using smaller and layer-specific low-rank perturbation matrices instead of a large and model-scale projection matrix, thus cutting memory and computational costs. Additionally, we introduce a lazy update strategy, generating perturbations periodically rather than iteratively, further reducing overhead. Besides, we also successfully apply SubZero to four popular LLM fine-tuning schemes, highlighting the compatibility of SubZero.
Secondly, we provide theoretical guarantees for SubZero. We first convert our gradient estimation into an equivalent formulation, highlighting the key differences between our approach and existing traditional ZO methods (Malladi et al., 2023), as well as random subspace ZO methods (Nozawa et al., 2024). Then, we prove that the gradient estimated by SubZero closely approximates the BP gradient, i.e., the ground-truth gradient, and enjoys significantly lower gradient variance than traditional ZO methods like MeZO. Furthermore, we establish the theoretical convergence of SubZero when combined with the SGD optimizer.
@@ 2 Related Work
Zeroth-Order Fine-Tuning. ZO optimizers utilize just two forward passes to estimate gradient without BP. Malladi et al. (2023) first used ZO optimization to fine-tune LLMs, significantly lowering the GPU hours and memory usage to levels similar to inference, which offers a considerable advantage over FO optimizers. They demonstrated that LLM fine-tuning benefits from a well-structured loss landscape by introducing suitable task-specific prompt templates. Convergence theories for ZO optimization have been elaborated in both convex (Nesterov & Spokoiny, 2017; Jamieson et al., 2012; Duchi et al., 2015) and non-convex settings (Liu et al., 2018; Ji et al., 2019). However, these convergence rates typically increase linearly with the number of trainable parameters (Nesterov & Spokoiny, 2017; Jamieson et al., 2012; Duchi et al., 2015; Liu et al., 2018; Ji et al., 2019).
Recently, more work in ZO has focused on improving the convergence rates and reducing gradient estimation variance for LLM fine-tuning. Increasing batch size can diminish noise in ZO gradient estimation (Gautam et al., 2024; Jiang et al., 2024). Perturbing a subset of model parameters also lowers gradient variance. This approach induces sparse parameter perturbations through random and sparse pruning masks (Liu et al., 2024) or block-coordinate perturbations (Zhang et al., 2024). Additionally, some approaches tried to reduce trainable parameters through PEFT (Malladi et al., 2023; Zhang et al., 2024) and tensorized adapters (Yang et al., 2024).
Random Subspace Optimization. To lessen dependence on dimensionality, some research utilizes random projections and low-dimensional perturbations in subspaces (Nozawa et al., 2024; Roberts & Royer, 2023; Kozak et al., 2021). However, these methods are hindered by the need to store a large projection matrix that increases with dimensionality, making it impractical for fine-tuning LLMs.
Memory-Efficient Fine-Tuning. Fine-tuning generally employs FO optimizers like SGD (Amari, 1993) or Adam (Kingma & Ba, 2015). Various approaches have been developed to reduce the memory cost of BP, such as sparsifying gradients (Sun et al., 2017), projecting gradients into a low-rank subspace (Zhao et al., 2024a), and quantizing optimizer states to lower bits (Dettmers et al., 2022b; Li et al., 2024). Additional methods to conserve activation and weight memory during forward and backward passes include gradient checkpointing (Chen et al., 2016), FlashAttention (Dao et al., 2022), QLoRA (Dettmers et al., 2024), and LLM.int8() (Dettmers et al., 2022a).
@@ 3 Preliminaries
In this section, we introduce the most popular ZO optimization approach and existing random subspace optimization methods.
Notations. We use a non-bold letter like $a$ and $A$ to denote a scalar, a boldfaced lower-case letter like $\bm{w}$ to denote a column vector, and a boldfaced upper-case letter such as $\bm{W}$ to denote a matrix. $\mathcal{N}(\bm{0},\bm{I})$ denotes a multivariate normal distribution with a zero mean vector and an identity covariance matrix. $\text{vec}(\bm{W})$ represents the vectorization of matrix $\bm{W}$, which transforms $\bm{W}$ into a column vector by stacking the columns of $\bm{W}$ vertically. $\bm{A}\otimes\bm{B}$ is the Kronecker product of matrices $\bm{A}$ and $\bm{B}$. $\mathbb{E}[{\bm{x}}]$ represents the expected value of a random variable ${\bm{x}}$. $\mathrm{Var}[{\bm{x}}]$ represents the variance of a random variable ${\bm{x}}$. The $\ell_{2}$-norm of a vector ${\bm{x}}$ is $\left\|{\bm{x}}\right\|=\sqrt{\sum_{i=1}^{n}{\bm{x}}_{i}^{2}}$. The spectral norm of a matrix ${\bm{A}}$ is $\left\|{\bm{A}}\right\|$. The Frobenius norm of a matrix $\bm{A}$ is $\|\bm{A}\|_{F}\!=\!\sqrt{\langle\bm{A},\bm{A}\rangle}$. $C^{s,p}_{L}(\mathcal{S})$ represents the class of $s$-th smooth and $p$-th $L$-smooth functions over the set $\mathcal{S}$. ${\rm bdiag}(\bm{A}_{1},\bm{A}_{2},\cdots,\bm{A}_{l})$ is a block diagonal matrix with diagonal blocks $\bm{A}_{1},\bm{A}_{2},\cdots,\bm{A}_{l}$.
We are interested in fine-tuning large LLMs (Ding et al., 2023). These models typically comprise multiple layers, with trainable parameter vectors represented as $\bm{w}\!=\!\left[\bm{w}_{1}^{\mathsf{T}},\bm{w}_{2}^{\mathsf{T}},\dots,\bm{w}_{l}^{\mathsf{T}}\right]^{\mathsf{T}}\!\in\!\mathbb{R}^{d}$, where $\bm{w}_{i}$ denotes the flattened parameter vector from the $i$-th layer and $d$ is the number of model parameters. Then training these models involves optimizing the following problem:
@@
Zeroth-Order Optimization. ZO optimization is BP-free and estimates gradients via random perturbations. A classical gradient estimator is the simultaneous perturbation stochastic approximation (SPSA) (Spall, 1992), which is defined as
@@
The SPSA in Eqn. (2) is an unbiased gradient estimator of the desired gradient $\nabla\mathbb{E}_{\bm{z}}[\mathcal{L}(\bm{w}+\varepsilon\bm{z})]$ (Nesterov & Spokoiny, 2017). It only requires two forward passes to estimate the gradient and eliminates the need for BP computation, resulting in substantial savings in computation cost and GPU memory usage. With this estimated gradient, it is easy to integrate with existing FO optimizers like SGD and develop corresponding ZO optimizers, such as ZO-SGD defined as:
@@
where $\eta^{t}>0$ is the learning rate at iteration $t$. To boost memory efficiency, MeZO (Malladi et al., 2023) implements ZO-SGD via in-place operations and employs a single random seed to facilitate efficient perturbation regeneration, significantly reducing memory overhead.
Random Subspace Optimization. Recent theoretical work (Nozawa et al., 2024; Roberts & Royer, 2023) has explored using low-dimensional perturbations in random subspaces to reduce gradient variances and enhance convergence rates. The key to random subspace methods is the generation of the perturbation vector $\tilde{\bm{z}}$ within a subspace spanned by $\bm{P}$:
@@
LLMs have a large model size, and thus their training and fine-tuning parameters can be very high-dimensional. This results in an excessively large matrix $\bm{P}$ which is $q$ times larger than the model size $d$ in full-parameter tuning (Aghajanyan et al., 2021) and is also large in other fine-tuning schemes e.g., LoRA (Hu et al., 2022). Consequently, this approach significantly increases memory requirements and computational complexity. Therefore, it is crucial to develop an efficient subspace construction strategy with minimal memory consumption for LLM fine-tuning.
@@ 4 Methodology
Here we first elaborate on our SubZero, a powerful ZO framework designed for LLM fine-tuning. Then we present how to integrate SubZero into four representative fine-tuning schemes.
@@ 4.1 Random Subspace Optimization for LLM Fine-Tuning
Our intuition is that exploring update directions in a low-dimensional subspace may result in a reduced variance of the estimated gradient compared to the estimation in the vanilla space as used in MeZO. Inspired by (Zhao et al., 2024a; Nozawa et al., 2024; Roberts & Royer, 2023), we propose the random Subspace Zeroth-order (SubZero) optimization framework tailored for LLM fine-tuning. This framework reduces gradient estimation variance, and minimizes the memory overhead associated with gradient estimation, such as the memory overhead caused by the projection matrix $\bm{P}$ in Eqn. (5) used in (Nozawa et al., 2024; Roberts & Royer, 2023).
@@
We propose a low-rank perturbation strategy for model parameter matrix of each layer, contrasting with previous random subspace methods that focus on the entire model’s parameters (Nozawa et al., 2024; Roberts & Royer, 2023). At each iteration, we generate a low-dimensional random matrix $\bm{Z}_{i}\in\mathbb{R}^{r\times r}$, where $r\ll\min\{m_{i},n_{i}\}$, and perform QR decomposition on two random matrices to create projection matrices $\bm{U}_{i}\in\mathbb{R}^{m_{i}\times r}$ and $\bm{V}_{i}\in\mathbb{R}^{n_{i}\times r}$ (see Algorithm 1). Both $\bm{U}_{i}$ and $\bm{V}_{i}$ are column-orthogonal matrices. Our experiments in Table 8 indicate that using Gaussian random projection matrices yields worse performance than using our designed column-orthogonal matrices. Then we combine these three matrices to yield a low-rank perturbation as follows:
@@
where $\tilde{\bm{Z}}_{i}$ is the perturbation matrix in a subspace spanned by $\bm{U}_{i}$ and $\bm{V}_{i}$, and $\bm{Z}_{i}$ represents the low-dimensional random perturbation matrix with entries sampled from $\mathcal{N}(0,1)$.
@@
In Sec. 5, we analyze the effectiveness of this new gradient estimation (8). Specifically, Theorem 5 proves the close distance between our gradient estimate (8) and the vanilla gradient computed by BP in FO methods, while Theorem 5 shows smaller variance and angle error of our gradient estimate in Eqn. (8) compared to the gradient estimate (2) in MeZO (Malladi et al., 2023). See more theoretical details in Sec. 5.
@@
Here we choose SGD as the default optimizer of SubZero. Theorem 5 in Sec. 5 guarantees the convergence of SubZero with SGD as basic optimizer and gives its convergence rate. The choice of FO optimizers is orthogonal to ZO optimization. However, some empirical work indicates that adaptive optimizers like Adam (Kingma & Ba, 2015) do not necessarily enhance convergence of ZO approaches during LLM fine-tuning (Zhang et al., 2024; Guo et al., 2024). Also, there are other ZO optimizers that utilize stochastic momentum (Jiang et al., 2024) and second-order information (Zhao et al., 2024b) to facilitate faster convergence. While SubZero can be applied with other FO and ZO optimizers, we leave a comprehensive evaluation of these approaches for future work.
Table 1: Comparison of memory cost between SubZero and representative optimizers in full-parameter tuning scheme with RoBERTa-large on SST-2. “Mem." represents the total GPU memory cost.
@@
We compare the memory overhead of SubZero with the existing random subspace method S-RGF (Nozawa et al., 2024) using identical experimental settings, including layer-wise perturbation and matching subspace dimension, with all methods utilizing the SGD optimizer. As shown in Table 1, S-RGF’s memory usage is roughly four times greater than SGD and 8.8 times that of MeZO (Malladi et al., 2023), while our SubZero’s memory usage is comparable to MeZO. See more experimental comparison on OPT-13B in Table 5 of Sec. 6.
Lazy Low-rank Subspace Update. According to Eqn. (9), at the $t$-th iteration, the gradient estimate of the parameter matrix in the $i$-th layer, ^$\widehat{\nabla}\mathcal{L}(\bm{W}_{i}^{t};\mathcal{B}^{t})$, lies within a subspace defined by the projection matrices $\bm{U}_{i}^{t}$ and $\bm{V}_{i}^{t}$. Specifically, $\bm{U}_{i}^{t}$ spans the column subspace, while $\bm{V}_{i}^{t}$ determines the row subspace, with both matrices generated iteratively, leading to extra computational overhead to LLM fine-tuning.
However, for LLM fine-tuning, enhancing the computational efficiency and the accuracy of gradient subspace approximation is crucial. An excessively short update interval for $\bm{U}_{i}$ and $\bm{V}_{i}$, such as generating them iteratively, can incur high computational costs and limit exploration of the gradient subspace they established. Conversely, a long interval may result in inaccuracies in subspace approximation and fail to capture the evolving nature of the gradient subspace. Inspired by Galore (Zhao et al., 2024a), we propose a lazy subspace update strategy that periodically regenerates the projection matrices $\bm{U}_{i}$ and $\bm{V}_{i}$. Specifically, these matrices are generated at the first iteration of every $T_{0}>1$ training iterations and remain unchanged for the subsequent $T_{0}-1$ iterations (see lines 4-7 in Algorithm 3). We utilize QR decomposition on two different random matrices for generating the column-orthogonal matrices $\bm{U}_{i}$ and $\bm{V}_{i}$, as summarized in Algorithm 1. This lazy subspace update strategy is both efficient and effective in all our experiments.

### v3

# Zeroth-Order Fine-Tuning of LLMs in Random Subspaces

@@ Abstract
Fine-tuning Large Language Models (LLMs) has proven effective for a variety of downstream tasks. However, as LLMs grow in size, the memory demands for backpropagation become increasingly prohibitive. Zeroth-order (ZO) optimization methods offer a memory-efficient alternative by using forward passes to estimate gradients, but the variance of gradient estimates typically scales linearly with the model’s parameter dimension—a significant issue for LLMs. In this paper, we propose the random Subspace Zeroth-order (SubZero) optimization to address the challenges posed by LLMs’ high dimensionality. We introduce a low-rank perturbation tailored for LLMs that significantly reduces memory consumption while improving performance. Additionally, we prove that our gradient estimation closely approximates the backpropagation gradient, exhibits lower variance than traditional ZO methods, and ensures convergence when combined with SGD. Experimental results show that SubZero enhances fine-tuning performance and achieves faster convergence compared to standard ZO approaches like MeZO across various language modeling tasks. Code is available at https://github.com/zimingyy/SubZero.
@@ 1 Introduction
Large Language Models (LLMs), such as the GPT and LLaMA series [64, 55], have recently demonstrated impressive capabilities in natural language processing tasks and beyond [52, 1]. These models utilize deep learning, particularly the transformer architecture [56], to learn complex patterns in language data. However, LLMs can struggle with specialized tasks that require domain-specific knowledge [49]. Fine-tuning presents an effective solution by slightly adjusting pre-trained LLMs with domain data, enabling them to adapt to specific tasks more effectively.
For fine-tuning, first-order (FO) optimizers, such as SGD [3] or Adam [30], are commonly used to achieve promising performance on domain datasets. However, as LLMs grow in size, FO optimizers demand increasingly memory consumption due to the gradient computations required by backpropagation (BP) [67]. Additionally, they are unable to directly handle non-differentiable objectives. To enhance memory efficiency, MeZO [40] first introduces the zeroth-order (ZO) optimizer to LLM fine-tuning without BP. It only requires forward passes and calculates gradient estimates using finite differences of training loss values, enabling it to directly handle non-differentiable objectives. Nevertheless, the variance of ZO gradient estimates scales linearly with the perturbation dimension, which corresponds to the number of model parameters. This can become extremely large in LLMs, leading to significant performance degradation compared to FO optimizers [20, 28, 39].
Reducing the variance of ZO gradient estimates generally results in faster convergence [62]. There are two main attempts to addressing the high variance of ZO gradient estimates. The first approach involves increasing batch size alongside training steps, which reduces gradient noise and variance in ZO gradient estimates [20, 28]. However, this leads to significant runtime and memory costs due to the large batch size in the later training stages. The second approach focuses on perturbing fewer parameters by employing sparse parameter perturbations, such as random and sparse pruning masks [39] and block-coordinate perturbations [65], or by reducing the number of trainable parameters through techniques like parameter-efficient fine-tuning (PEFT) [40, 65] and tensorized adapters [60]. Recent theoretical advancements have proposed using random projections to lessen the dimensionality dependence in ZO optimizers [43, 46, 31] by applying low-dimensional perturbations in random subspaces. Nonetheless, a major drawback of this approach is the need to store a huge projection matrix that scales with model parameter dimensionality, making it impractical for fine-tuning large LLMs.
@@
Figure 1: Visualization of cosine similarity $\mathbb{E}\left[{\texttt{cosine}(\bm{g},\hat{\bm{g}})}\right]$, relative variance $\mathrm{Var}\bigl{[}\left\|\hat{\bm{g}}\right\|\bigr{]}/\left\|\bm{g}\right\|^{2}$, training loss, and peak total GPU memory cost with OPT-1.3B on SST-2 in the prompt tuning scheme. All three methods utilize a batch size of 16 and run for 20K steps. Here, $\hat{\bm{g}}$ represents the gradient estimated by MeZO or our SubZero, and $\bm{g}$ denotes the expected gradient $\mathbb{E}[\hat{\bm{g}}]$. Theorem 5 (b) ensures that SubZero maintains a small distance between $\bm{g}$ and the BP gradient in a subspace. (a) and (b) demonstrate that SubZero’s estimated gradient $\hat{\bm{g}}$ has lower angle error and variance than MeZO. (c) and (d) indicate that SubZero enhances convergence speed with minimal extra memory usage.
@@
Firstly, we propose a layer-wise low-rank perturbation approach for gradient estimation, specifically designed for fine-tuning LLMs. In each layer, we generate a low-rank perturbation matrix by combining two column-orthogonal matrices with a Gaussian random matrix, which is then used for gradient estimation. Unlike traditional ZO methods like MeZO [40] which apply non-low-rank perturbations to the entire model, our approach significantly reduces the variance of gradient estimates and the angle error between the estimated gradient and its expectation, as respectively shown in Fig. 1 (a) and (b). SubZero also improves upon random subspace ZO methods like S-RGF [43] by using smaller and layer-specific low-rank perturbation matrices instead of a large and model-scale projection matrix, thus cutting memory and computational costs. Additionally, we introduce a lazy update strategy, generating perturbations periodically rather than iteratively, further reducing overhead. Besides, we successfully apply SubZero to four popular LLM fine-tuning schemes, highlighting the compatibility of SubZero.
Secondly, we provide theoretical guarantees for SubZero. We first convert our gradient estimation into an equivalent formulation, highlighting the key differences between our approach and existing traditional ZO methods [40], as well as random subspace ZO methods [43]. Then, we prove that the gradient estimated by SubZero closely approximates the BP gradient, i.e., the ground-truth gradient, and enjoys significantly lower gradient variance than traditional ZO methods like MeZO. Furthermore, we establish the theoretical convergence of SubZero when combined with the SGD optimizer.
@@ 2 Related Work
Zeroth-Order Fine-Tuning. ZO optimizers utilize just two forward passes to estimate gradient without BP. Malladi et al. [40] first used ZO optimization to fine-tune LLMs, significantly lowering the GPU hours and memory usage to levels similar to inference, which offers a considerable advantage over FO optimizers. They demonstrated that LLM fine-tuning benefits from a well-structured loss landscape by introducing suitable task-specific prompt templates. Convergence theories for ZO optimization have been elaborated in both convex [41, 25, 19] and non-convex settings [37, 26]. However, these convergence rates typically increase linearly with the number of trainable parameters [41, 25, 19, 37, 26].
Recently, more work in ZO has focused on improving the convergence rates and reducing gradient estimation variance for LLM fine-tuning. Increasing batch size can diminish noise in ZO gradient estimation [20, 28]. Perturbing a subset of model parameters also lowers gradient variance. This approach induces sparse parameter perturbations through random and sparse pruning masks [39] or block-coordinate perturbations [65]. Additionally, some approaches tried to reduce trainable parameters through PEFT [40, 65] and tensorized adapters [60].
Random Subspace Optimization. To lessen dependence on dimensionality, some research utilizes random projections and low-dimensional perturbations in subspaces [43, 46, 31]. However, these methods are hindered by the need to store a large projection matrix that increases with dimensionality, making it impractical for fine-tuning LLMs.
Memory-Efficient Fine-Tuning. Fine-tuning generally employs FO optimizers like SGD [3] or Adam [30]. Various approaches have been developed to reduce the memory cost of BP, such as LoRA [24, 58], gradient sparsification [54], low-rank gradient projection [67], and optimizer state quantization [15, 35]. Additional methods to conserve activation and weight memory during forward and backward passes include gradient checkpointing [8], FlashAttention [12], QLoRA [16], and LLM.int8() [14].
@@ 3 Preliminaries
Here we introduce the most popular ZO optimization approach and existing random subspace optimization methods.
Notations. Let non-bold letter like $a$ and $A$ denote a scalar, a boldfaced lower-case letter like $\bm{w}$ denote a column vector, and a boldfaced upper-case letter like $\bm{W}$ denote a matrix. $\mathcal{N}(\bm{0},\bm{I})$ is a multivariate normal distribution with a zero mean vector and an identity covariance matrix. $\text{vec}(\bm{W})$ denotes the vectorization of matrix $\bm{W}$ which reshapes $\bm{W}$ into a column vector by stacking the columns of $\bm{W}$ vertically. $\bm{A}\otimes\bm{B}$ is the Kronecker product of matrices $\bm{A}$ and $\bm{B}$. $\mathbb{E}[{\bm{x}}]$ is the expected value of a random variable ${\bm{x}}$. $\mathrm{Var}[{\bm{x}}]$ is the variance of a random variable ${\bm{x}}$. The $\ell_{2}$-norm of a vector ${\bm{x}}$ is $\left\|{\bm{x}}\right\|=\sqrt{\sum_{i=1}^{n}{\bm{x}}_{i}^{2}}$. The spectral norm of a matrix ${\bm{A}}$ is $\left\|{\bm{A}}\right\|$. The Frobenius norm of a matrix $\bm{A}$ is $\|\bm{A}\|_{F}\!=\!\sqrt{\langle\bm{A},\bm{A}\rangle}$. $C^{s,p}_{L}(\mathcal{S})$ denotes the class of $s$-th smooth and $p$-th $L$-smooth functions over the set $\mathcal{S}$. ${\rm bdiag}(\bm{A}_{1},\bm{A}_{2},\cdots,\bm{A}_{l})$ is a block diagonal matrix with diagonal blocks $\bm{A}_{1},\bm{A}_{2},\cdots,\bm{A}_{l}$.
We are interested in fine-tuning large LLMs [17]. These models typically comprise multiple layers, with trainable parameter vectors represented as $\bm{w}\!=\!\left[\bm{w}_{1}^{\mathsf{T}},\bm{w}_{2}^{\mathsf{T}},\dots,\bm{w}_{l}^{\mathsf{T}}\right]^{\mathsf{T}}\!\in\!\mathbb{R}^{d}$, where $\bm{w}_{i}$ denotes the flattened parameter vector from the $i$-th layer and $d$ is model parameter dimension. Then training these models involves optimizing the problem:
@@
Zeroth-Order Optimization. ZO optimization is BP-free and estimates gradients via random perturbations. A classical gradient estimator is the simultaneous perturbation stochastic approximation (SPSA) [53], which is defined as
@@
The SPSA in Eqn. (2) is an unbiased gradient estimator of the desired gradient $\nabla\mathbb{E}_{\bm{z}}[\mathcal{L}(\bm{w}+\varepsilon\bm{z})]$ [41]. It only requires two forward passes to estimate the gradient and eliminates BP computation, greatly reducing computation cost and GPU memory. With this estimated gradient, one can integrate with existing FO optimizers like SGD to develop corresponding ZO optimizers, e.g., ZO-SGD defined as:
@@
where $\eta^{t}$ is the learning rate at step $t$. In practice, MeZO [40] implements ZO-SGD via in-place operations and uses a single random seed to facilitate efficient perturbation regeneration, greatly reducing memory overhead.
Random Subspace Optimization. Recent theoretical work [43, 46] has explored using low-dimensional perturbations in random subspaces to reduce gradient variances and hence enhance convergence rates. The key to random subspace methods is the generation of the perturbation vector $\tilde{\bm{z}}$ within a subspace spanned by $\bm{P}$:
@@
LLMs have a large model size, and thus their training and fine-tuning parameters can be very high-dimensional. This results in an excessively large matrix $\bm{P}$ which is $q$ times larger than the model size $d$ in full-parameter tuning [2] and is also large in other fine-tuning schemes e.g., LoRA [24]. Consequently, this approach significantly increases memory requirements and computational complexity. Therefore, it is crucial to develop an efficient subspace construction strategy with minimal memory consumption for LLM fine-tuning.
@@ 4 Methodology
Here we first elaborate on our SubZero, a powerful ZO framework for LLM fine-tuning. Then we present how to integrate SubZero into four representative fine-tuning schemes.
@@ 4.1 Random Subspace Optimization for LLM Fine-Tuning
Our intuition is that exploring update directions in a low-dimensional subspace may result in a reduced variance of the estimated gradient [43, 46] compared to the estimation in the vanilla space as used in MeZO. Moreover, recent work indicates that BP gradients in LLM fine-tuning rapidly converge to a small subspace [66, 40, 67, 23]. Accordingly, we propose the random Subspace Zeroth-order (SubZero) optimization framework tailored for LLM fine-tuning. This framework reduces gradient estimation variance, and minimizes the memory overhead associated with gradient estimation, such as the memory overhead caused by the projection matrix $\bm{P}$ in Eqn. (5) used in [43, 46].
@@
We propose a low-rank perturbation strategy for model parameter matrix of each layer, contrasting with previous random subspace methods that focus on the entire model’s parameters [43, 46]. At each step, we generate a low-dimensional random matrix $\bm{Z}_{i}\in\mathbb{R}^{r\times r}$, where $r\ll\min\{m_{i},n_{i}\}$, and perform QR decomposition on two Gaussian random matrices with entries sampled from $\mathcal{N}(0,1)$ to create projection matrices $\bm{U}_{i}\in\mathbb{R}^{m_{i}\times r}$ and $\bm{V}_{i}\in\mathbb{R}^{n_{i}\times r}$ (see Algorithm 1). Both $\bm{U}_{i}$ and $\bm{V}_{i}$ are column-orthogonal matrices. Our experiments in Table 5 indicate that directly using Gaussian random projection matrices yields worse performance than using our designed column-orthogonal matrices. Then we combine these three matrices to yield a low-rank perturbation as follows:
@@
where $\tilde{\bm{Z}}_{i}$ is the perturbation matrix in a subspace spanned by $\bm{U}_{i}$ and $\bm{V}_{i}$, and $\bm{Z}_{i}$ represents the low-dimensional random perturbation matrix with entries sampled from $\mathcal{N}(0,1)$. The projection matrices for SubZero can only be derived using the QR decomposition of the weights, activations, ZO gradients, or random matrix, without increasing memory overhead. As shown in Table 2, our proposed random matrix achieves the best performance. While the weight and ZO gradient matrices are feasible, their performance drops significantly. The activation matrix is ineffective due to its batch size dependency, requiring more sophisticated handling. Detailed experimental setups are in Appendix 8.4.
@@
In Sec. 5, we analyze the effectiveness of this new gradient estimation (8). Specifically, Theorem 5 proves the close distance between our gradient estimate (8) and the vanilla gradient computed by BP in FO methods, while Theorem 5 shows smaller variance and angle error of our gradient estimate in Eqn. (8) compared to the gradient estimate (2) in MeZO [40]. See more theoretical details in Sec. 5.
@@
Here we choose SGD as the default optimizer of SubZero. Theorem 5 in Sec. 5 guarantees the convergence of SubZero with SGD as basic optimizer and gives its convergence rate. The choice of FO optimizers is orthogonal to ZO optimization. Also, some empirical work indicates that adaptive optimizers like Adam [30] do not necessarily enhance convergence of ZO approaches during LLM fine-tuning [65, 22]. So the combination of SubZero and Adam is included in Appendix 8.1 due to the limited space. We apply the primitive ZO approach. There are other ZO optimizers that utilize stochastic momentum [28] and second-order information [68] to facilitate faster convergence. While SubZero can be adapted to these ZO optimizers, we leave a comprehensive evaluation of these approaches for future work.
Table 1: Projection matrix generation for SubZero in full-parameter tuning with OPT-1.3B and LLaMA2-7B on SST-2 and OPT-13B on RTE.
Matrix 1.3B 7B 13B
Weight 91.5 91.7 65.3
Activation 51.5 52.9 53.1
ZO Gradient 89.6 92.0 67.5
Random 93.4 94.5 74.0
Table 2: Memory cost in full-parameter tuning with RoBERTa-large on SST-2.
@@
We compare the memory overhead of SubZero with the existing random subspace method S-RGF [43] using identical experimental settings, including layer-wise perturbation and matching subspace dimension, with all methods utilizing the SGD optimizer. As shown in Table 2, S-RGF’s memory usage is roughly four times greater than SGD and 8.8 times that of MeZO [40], while our SubZero’s memory usage is comparable to MeZO. See more experimental comparisons on OPT-13B in Table 10 of Appendix 8.1.
Lazy Low-rank Subspace Update. According to Eqn. (9), at the $t$-th step, the gradient estimate of the parameter matrix in the $i$-th layer, ^$\widehat{\nabla}\mathcal{L}(\bm{W}_{i}^{t};\mathcal{B}^{t})$, lies within a subspace defined by the projection matrices $\bm{U}_{i}^{t}$ and $\bm{V}_{i}^{t}$. Specifically, $\bm{U}_{i}^{t}$ spans the column subspace, while $\bm{V}_{i}^{t}$ determines the row subspace, with both matrices generated iteratively, leading to extra computational overhead to LLM fine-tuning.
However, for LLM fine-tuning, enhancing the computational efficiency and the accuracy of gradient subspace approximation is crucial. An excessively short update interval for $\bm{U}_{i}$ and $\bm{V}_{i}$, such as generating them iteratively, can incur high computational costs and limit exploration of the gradient subspace they established. Conversely, a long interval may result in inaccuracies in subspace approximation and fail to capture the evolving nature of the gradient subspace. Accordingly, we propose a lazy subspace update strategy that periodically regenerates the projection matrices $\bm{U}_{i}$ and $\bm{V}_{i}$. Specifically, these matrices are generated at the first step of every $F>1$ training steps and remain unchanged for the subsequent $F-1$ steps (see lines 4-7 in Algorithm 3). We utilize QR decomposition on two different random matrices for generating the column-orthogonal matrices $\bm{U}_{i}$ and $\bm{V}_{i}$, as summarized in Algorithm 1. This lazy subspace update strategy is both efficient and effective in all our experiments.